OpenStudy (anonymous):

The base of a square pyramid has a side length of 7 inches. The height of the pyramid is 19 inches. What is the approximate volume of the pyramid? A. 133 in3 B. 177 in3 C. 310 in3 D. 931 in3

OpenStudy (texaschic101):

V = a^2 * h/3 a = 7 h = 19 now sub V = 7^2 * 19/3 V = 49 * 19/3 V = 931/3 V = 310.33....rounds to 310 in^3